OpenStudy (majesty69):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can do this 10 times . Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can receive.

OpenStudy (majesty69):

I need help starting this problem please :)

OpenStudy (anonymous):

What is the total test score possible? 100?

OpenStudy (majesty69):

Yes

OpenStudy (anonymous):

So first you need to make an equation. Do you know what form it should be in?

OpenStudy (majesty69):

A geometric series Sn=a1-a1(r)^n / 1-r

OpenStudy (anonymous):

Sorry I do not know how that posted

OpenStudy (majesty69):

Oh haha it's okay! Was I right?

OpenStudy (anonymous):

Well you would gain 30% of you test score back every time. I would use an equation such as y=a(b)^x where x is number of entries, a is first test score, and b is max increase possible

OpenStudy (majesty69):

But it has to be a in geometric series form is the thing. So wouldnt it be S10 = 65 - (65)3/10^10

OpenStudy (majesty69):

I think I got the answer... after you plug it in you get 64.999/(7/10) = 92.85 So she can get a maximum of 92.9 % after ten times.

OpenStudy (anonymous):

that looks about right